OpenStudy (anonymous):

integrate x(x-1)^(1/2) lower bound = 1 upper bound = 2 I know the answer but when I try and do it out I keep getting the wrong one, any help?

myininaya (myininaya):

\[\int\limits_{1}^{2}x \sqrt{x-1} dx\] Did you do u=x-1 => du =dx Also u=x-1 => u+1=x And if x=1 then u=1-1=0 And if x=2 then u=2-1=1 So we have \[\int\limits_{0}^{1} (u+1) \sqrt{u} du\]

myininaya (myininaya):

Did you get this far?

myininaya (myininaya):

oh and i like robots :)

OpenStudy (shayaan_mustafa):

myininaya. i need your help. I messaged you. but you didn't reply me. :(

OpenStudy (anonymous):

i cant see what you got because it says math processing answer

myininaya (myininaya):

Try refreshing...

myininaya (myininaya):

@imalittletoyrobotfromthefuture did that work?

OpenStudy (anonymous):

yea it worked, i have that now. I didnt know what to do because I didnt recognize that x = u + 1

myininaya (myininaya):

Well if u=x-1 then u+1=x (I just added 1 on both sides for u=x-1)

OpenStudy (anonymous):

yes... haha i got that, what i had before though was: x\[\int\limits_{1}^{2}x \sqrt{u}du\]

myininaya (myininaya):

But your integral needs to be in terms of one variable and when you make the switch from x to u your limits also need to switch from x to u

OpenStudy (anonymous):

yes, thats why i was having trouble, i didnt recognize the fact that I could put x in terms of u, now that I have that though I'm unsure of what to do next

myininaya (myininaya):

\[\int\limits\limits_{0}^{1} (u+1) \sqrt{u} du \] So this is what I wrote above

myininaya (myininaya):

Do you understand exactly how I got it?

OpenStudy (anonymous):

yes, what next though? this?: \[\int\limits_{0}^{1}((u^2/2)+u)((2u^3/2)/3)\]

myininaya (myininaya):

?

OpenStudy (anonymous):

isn't that the anti-derivitave of the equation \[\int\limits_{0}^{1}(u+1)\sqrt{u}du\]?

myininaya (myininaya):

\[\int\limits_{0}^{1}(u+1) \sqrt{u} du=\int\limits_{0}^{1}(u+1)u^\frac{1}{2} du=\int\limits_{0}^{1}(u \cdot u^\frac{1}{2}+1 \cdot u^\frac{1}{2}) du\] \[=\int\limits_{0}^{1}(u^\frac{3}{2}+u^\frac{1}{2}) du\]

myininaya (myininaya):

Now we integrate...

myininaya (myininaya):

\[[\frac{u^{\frac{3}{2}+1}}{\frac{3}{2}+1}+\frac{u^{\frac{1}{2}+1}}{\frac{1}{2}+1}]_0^1\]

myininaya (myininaya):

\[=[\frac{1^{\frac{3}{2}+1}}{\frac{3}{2}+1}+\frac{1^{\frac{1}{2}+1}}{\frac{1}{2}+1}]-[\frac{0^{\frac{3}{2}+1}}{\frac{3}{2}+1}+\frac{0^{\frac{1}{2}+1}}{\frac{1}{2}+1}]\]

myininaya (myininaya):

robot you have questions?

OpenStudy (anonymous):

16/15 :) thanks for the explanation, very helpful and yes, robots are sick. check this out http://www.cnn.com/2012/03/04/opinion/ted-kumar-flying-robots/index.html?iref=obnetwork

myininaya (myininaya):

lol